# Gradient descent induces alignment between weights and the empirical NTK for deep non-linear networks

### v1

@@ 2.2 Alignment decomposition
This alignment holds trivially and exactly if $K^{(\ell)}$ is the identity. However, the correlation can be high in trained networks even with non-trivial $K^{(\ell)}$. For example, in the chain monomial task (Figure 1), $K^{(0)}$ is far from identity (standard deviation of its eigenvalues is $5.9$ times its average eigenvalue), but the NFA correlation is $0.93$ at the end of training. We also note that if $K^{(\ell)}$ is independent of $W^{(\ell)}$, the alignment is lower than in trained networks; in the same example, replacing $K^{(0)}$ with a matrix $Q$ with equal spectrum but random eigenvectors greatly reduces the correlation to $0.53$ and qualitatively disrupts the structure relative to the NFM (Figure 1, rightmost column). We show the same result for the CelebA dataset (see Appendix K). Therefore, the NFA has to do with the alignment of the left eigenvectors of $W^{(\ell)}$ with $K^{(\ell)}$ in addition to spectral considerations.
The $k_{\ell}\times k_{\ell}$ matrix $K^{(\ell)}$ is the feature covariance of the pre-activation tangent kernel (PTK) $\mathcal{K}^{(\ell)}$ which we define as:
$\mathcal{K}^{(\ell)}(x,z)\equiv\frac{\partial f(x)}{\partial h_{\ell}}\cdot\frac{\partial f(z)}{\partial h_{\ell}}$ (2)
The PTK is related to the layerwise empirical NTK (ENTK) $\hat{\theta}_{\ell}$ by the following identity:
Figure 2: Uncentered and centered NFA correlations for a two hidden layer MLP trained on data drawn from a standard Normal distribution on the chain monomial task of rank $r=5$. Only, the initialization scale of the first layer weights is varied, while $s_{\ell}=1$ for $\ell>0$. The top row shows the values for layer 1, while the bottom row are the values for layer 2. Train (test) losses are scaled by the maximum train (test) loss achieved so that they are between 00 and $1$. Here width is $256$, the input dimension is $32$, and the dataset contains $256$ points.
Proposition 3 (Pre-activation to neural tangent identity).
Consider a depth-$L$ neural network $f(x)$ with inputs $X_{\ell}\in\mathbb{R}^{n\times k_{\ell}}$ to weight matrices $W^{(\ell)}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$ for $\ell\in[L]$. Consider the empirical NTK where gradients are taken only with respect to $W^{(\ell)}$ evaluated between two points $x$ and $z$ denoted by $\hat{\Theta}_{\ell}(x_{0},z_{0})=\langle\frac{\partial f(x_{0})}{\partial W^{(\ell)}},\frac{\partial f(z_{0})}{\partial W^{(\ell)}}\rangle$. Then for all $\ell$, Θ^subscriptnormal-^normal-$\hat{\Theta}_{\ell}$ satisfies,
Therefore, the NFA is a statement about the alignment of the weight matrices with the PTK feature covariance - which itself is a factor of the layerwise NTK.
2.3 Centering the NFA isolates weight-PTK alignment
We showed that the NFA is equivalent to PTK-weight alignment (Proposition 2), and that it emerges during training (Figure 2 and Appendix G). We now ask: is the development of the NFA due to alignment of the weight matrices to the current PTK, or the alignment of the PTK to the current weights?
In practice, both effects matter, but numerical evidence suggests that changes in the PTK don’t drive the early dynamics of the NFA (Appendix B). Instead, we focus on the alignment of the weights to the PTK at early times. We can measure this alignment using an alternative, centered version of the NFA:
Ansatz 4 (Centered Neural Feature Ansatz).
Let $W_{0}^{(\ell)}$, $W^{(\ell)}$, and $K^{(\ell)}$ be the initial weight matrix, trained weight matrix, and interior feature matrix, respectively, at layer $\ell$, and let $\bar{W}^{(\ell)}\equiv W^{(\ell)}-W^{(\ell)}_{0}$. The centered neural feature ansatz (C-NFA) postulates that the following quantities will remain proportional to one another,
$(\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)}\propto(\bar{W}^{(\ell)})^{\top}K^{(\ell)}\bar{W}^{(\ell)}~{}.$
We refer to the centered NFA as the C-NFA, and the original, uncentered NFA as the UC-NFA to distinguish them.
Since $\bar{W}=0$ at initialization, the early dynamics of the C-NFA are dominated by $W$ aligning with the initial $K$:
Proposition 5 (Centered NFA dynamics).
At initialization, $\bar{W}^{\top}\bar{W}=\bar{W}^{\top}K\bar{W}=0$, and,
$\begin{split}\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=0,\quad&\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=2\dot{W}^{\top}\dot{W},\\ \frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0,\quad&\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})=2\dot{W}^{\top}K\dot{W}~{}.\end{split}$
The first non-zero derivatives of the two quantities in the C-NFA give us the change in the gram matrix of $W$, as well as the change in the AGOP for fixed PTK $K$. This makes it a useful object to study weight-PTK alignment.
We note that the C-NFA correlation remains high throughout training across initialization scale (Figure 2, second column, and Appendix G), but is especially high and invariant to scale early on in training. This is in contrast to the UC-NFA, whose final value can be extremely sensitive to the initialization scale (e.g. second row, Figure 2).
In the remainder of this work, we first analyze the C-NFA in early training and show that gradient based training drives the C-NFA, with a strength that depends on model and datasets. (Section 3). We then discuss initializations and optimizers which can increase the strength of the C-NFA and induce larger correlation between $F$ and $\bar{G}$ (Section 4).
3 Theoretically predicting the centered NFA
3.1 Early time C-NFA dynamics
To understand the dynamics of the weights aligning to the PTK, we compute the time derivatives of the C-NFA at initialization (Appendix A):
Proposition 6 (Centered NFA decomposition).
For a fully-connected neural network at initialization, whose weights are trained with gradient descent on a loss function $\mathcal{L}$,
$\dot{W}^{\top}\dot{W}=X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X,\quad\dot{W}^{\top}K\dot{W}=X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}$ (3)
@@ 3.1 Early time C-NFA dynamics
where $\dot{\mathcal{L}}$ is the $n\times n$ diagonal matrix of logit derivatives
˙$\dot{\mathcal{L}}\equiv{{\rm diag}\left(\frac{\partial\mathcal{L}}{\partial f}\right)}$
We immediately see from these equations how gradient-based training can drive the NFA. Even though the NFA doesn’t explicitly involve the loss, the gradient descent dynamics of the NFM and AGOP depend on the labels in a similar way, and differ by a factor of the kernel $\mathcal{K}$ similar to how the AGOP and NFM differ by a factor of $K$.
@@
In order to understand the correlation more quantitatively, we will focus on the case of mean-squared error (MSE) loss in the rest of this work. Here $\dot{\mathcal{L}}$ corresponds to the diagonal matrix of the residuals $y-f(x)$. If the outputs of the network is 00 on the training data, then ˙$\dot{\mathcal{L}}=Y\equiv{\rm diag}\left(y\right)$, the labels themselves. In that case, the correlation of the time derivatives at initialization is a function of the input-label covariances as well as the PTK matrix:
@@
$\begin{split}&\mathbb{E}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right]=\\ &\text{tr}\left(X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YXX^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX\right)\\ &+\text{tr}\left(\text{Cov}\left(X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\right)\right)\end{split}$ (5)
A similar decomposition exist for the denominator terms.
The simplest limit is the NTK regime where the width tends to infinity with fixed number of data points and input dimension. Here the PTK matrix will approach its expectation, and the second, covariance term above will tend to 0. This suggests that the first (mean) term encodes much of the interesting phenomenology near initialization for large networks. We use this fact to design a dataset that interpolates between small and large initial derivatives of the C-NFA, even in the non-NTK regime, which will help us demonstrate the validity of our random matrix theory approach.
We focus most of our random matrix theory analysis on a one-hidden layer network, as the calculations quickly become complicated with depth. However, we also show that the mean term can be used to approximately predict the NFA value for Gaussian data with different spectral decay rates for more complicated networks (see Appendix C).
3.2 Exact predictions with one hidden layer, quadratic activation
@@
To illustrate our analysis, we derive here the numerator of Equation (4). Let $M^{(4)}_{X|Y}=(X^{\top}YX)^{2}$, and $M^{(2)}_{X}=X^{\top}X$, and $F_{a}=W^{\top}{\rm diag}\left(a^{2}\right)W$. Then,
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)$
$=\text{tr}\left(M^{(4)}_{X|Y}F_{a}M^{(4)}_{X|Y}F_{a}M^{(2)}_{X}F_{a}\right)~{}.$
@@
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)$
$=\text{tr}\left(M^{(4)}_{X|Y}\right)^{2}\text{tr}\left(F_{a}^{3}\right)+\text{tr}\left(F_{a}\right)\text{tr}\left(F_{a}^{2}\right)\text{tr}\left((\bar{M}^{(4)}_{X|Y})^{2}\right)~{}.$
@@
To numerically explore the validity of the random matrix theory calculations, we need a way to generate datasets with different values of $\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)$. We construct a random dataset called the alignment reversing dataset, parameterized by a balance parameter $\gamma\in(0,1]$ to adversarially disrupt the NFA near initialization in the regime that width $k$, input dimension $d$, and dataset size $n$ are all equal ($n=k=d=1024$). By Proposition 12, for the aforementioned neural architecture, the expected second derivative of the centered NFM satisfies, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YX=(X^{\top}YX)^{2}$, while the expected second derivative of the centered AGOP, $\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX$, has an additional component $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$. Our construction exploits this difference in that $X^{\top}X$ becomes adversarially unaligned to $X^{\top}YX$ as the balance parameter decreases. In our experiment, we sample multiple random datasets with this construction and compute the predicted and observed correlation of the second derivatives of the centered NFA at initialization.
@@
The construction exploits that we can manipulate $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$ freely of the NFM using a certain choice of $Y$ (see Appendix F for details of the construction). We design the dataset such that this AGOP-unique term is close to identity, while the NFM second derivative has many large off-diagonal entries, leading to low correlation between the second derivatives of the NFM and AGOP.
@@
We observe in Figure 3 that the centered NFA correlations predicted with random matrix theory closely match the observed values, across individual four random seeds and for the average of the correlation across them. Crucially, a single neural network is used across the datasets, confirming the validity of the self-averaging assumption. The variation in the plot across seeds come from randomness in the sample of the data, which cause deviations from the adversarial construction.
4 Increasing the centered contribution to the uncentered NFA improves feature learning

### v2

@@ 2.2 Alignment decomposition
This alignment holds trivially and exactly if $K^{(\ell)}$ is the identity. However, the correlation can be high in trained networks where $K^{(\ell)}$ is non-trivial. For example, in the chain monomial task (Figure 1), $K^{(0)}$ is far from identity (standard deviation of its eigenvalues is $5.9$ times its average eigenvalue), but the NFA correlation is $0.93$ at the end of training. We also note that if $K^{(\ell)}$ is independent of $W^{(\ell)}$, the alignment is lower than in trained networks; in the same example, replacing $K^{(0)}$ with a matrix $Q$ with equal spectrum but random eigenvectors greatly reduces the correlation to $0.53$ and qualitatively disrupts the structure relative to the NFM (Figure 1, rightmost column). We show the same result for the CelebA dataset (see Appendix L). Therefore, the NFA is a consequence of alignment between the left eigenvectors of $W^{(\ell)}$ and $K^{(\ell)}$ in addition to spectral considerations.
Note that $K^{(\ell)}$ itself is the AGOP of the neural network $f$ with respect to the pre-activations at layer $\ell$, i.e. $K^{(\ell)}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial h_{\ell}}^{\top}$. For convolutional and attention layers, the $K^{(\ell)}$ are computed by additionally averaged over all patches and token positions in the input, respectively (see Radhakrishnan et al. [21], Beaglehole et al. [22] for the formulation of the NFA in these architectures).
Figure 2: Uncentered and centered neural feature correlations across (A,B) fully-connected, (C) convolutional, and (D) attention layers with large initialization scale. (A,C,D) show trajectories of C/UC-NFC over training. (B) shows NFC values across all layers of an MLP with five hidden layers, averaged over CIFAR-10, CIFAR-100, SVHN, MNIST, GTSRB, and STL-10 datasets. (A-C) are additionally averaged over three random seeds. Each row of (D) is an attention block (ordered from first to last in the GPT model), while the columns show correlations for query, key, and value layers, respectively.
3 Centering the NFC isolates weight-PTK alignment
We showed that the neural feature ansatz is equivalent to PTK-weight alignment (Proposition 2). We now ask: is the increase in the NFC due to alignment of the weight matrices to the current PTK, or the alignment of the PTK to the current weights? In practice, both effects matter, but numerical evidence suggests that changes in the PTK do not drive the early dynamics of the NFC (Appendix C). Instead, we observe that the alignment between the weights and the PTK feature covariance is driven by a centered NFC that captures alignment between the parameter changes and the PTK. We then show this centered NFC can hold robustly in settings where the NFC holds with correlation less than 1, such as early in training and/or with large initialization. Finally, we establish analytically through the centered NFC how the NFA emerges as a structural property of gradient descent.
We begin by considering a decomposition of the NFM and AGOP into parts that depend on initialization and a part that depends on the changes in the weight matrix. Let $W^{(\ell)}_{t}$ and $K^{(\ell)}_{t}$ be the weight matrix and PTK feature covariance for layer $\ell$ at time $t$. We can write the NFM and AGOP in terms of the initial weights $W^{(\ell)}_{0}$ and the change in weights $\bar{W}^{(\ell)}_{t}\equiv W^{(\ell)}_{t}-W^{(\ell)}_{0}$ as follows:
$\begin{split}F{}=W_{t}^{\top}W_{t}&=\bar{W}_{t}^{\top}\bar{W}_{t}+W_{0}^{\top}\bar{W}_{t}+\bar{W}_{t}^{\top}W_{0}+W_{0}^{\top}W_{0},\\ \bar{G}{}=W_{t}^{\top}K_{t}W_{t}&=\bar{W}_{t}^{\top}K_{t}\bar{W}_{t}+W_{0}^{\top}K_{t}\bar{W}_{t}+\bar{W}_{t}^{\top}K_{t}W_{0}+W_{0}^{\top}K_{t}W_{0}\end{split}$ (3)
where we omitted $\ell$ from all terms for ease of notation. The first term in each decomposition isolates the gram matrix of the changes in the weight matrix. We call $\bar{W}^{\top}\bar{W}$ the centered NFM and $\bar{W}^{\top}K\bar{W}$ the centered AGOP. The centered AGOP in particular measures the alignment of weight updates with the current PTK feature covariance. Both terms are 00 at initialization, and if the weight matrices change significantly (that is, if $||\bar{W}||\gg||W_{0}||$), both the NFM and AGOP are dominated by the centered terms.
The decomposition in Equation 3 suggests that updates which drive correlation of the centered NFM and centered AGOP can also drive the overall value of the original NFC. Inspired by this observation, we define the centered NFC (C-NFC) as,
$\rho\left((\bar{W}^{(\ell)})^{\top}\bar{W}^{(\ell)},(\bar{W}^{(\ell)})^{\top}K^{(\ell)}\bar{W}^{(\ell)}\right)~{}.$
For the remainder of the paper, we will refer to the original NFC as the uncentered NFC (UC-NFC) to avoid ambiguity.
The centered NFC is consistently higher than the uncentered NFC across training times, architectures, and datasets Figure 2. This especially holds at early times and in deeper layers of a network (Figure 2C and D, VGG-11 on CIFAR-10 and the GPT model on Shakespeare respectively). For MLPs, we conduct a broader set of experiments and verify that the trends hold on a wide range of vision datasets. We also note that the C-NFC at early times is relatively robust to the initialization statistics of the weight matrix $W^{(\ell)}$ unlike the UC-NFC (for additional experiments, see Appendix G). We will return to this point in Section 5. These experiments suggest the C-NFC is responsible for improvements in the uncentered correlation.
High C-NFC values can drive increases in the UC-NFC as long as the centered NFM and centered AGOP are increasing in magnitude during training. We can confirm that the weights move significantly from initialization, which drives the contribution from the centered NFM and centered AGOP (Figures 14, and 15 in Appendix L). The increased importance of the C-NFC leads the UC-NFC to converge to the C-NFC at late times in most of our experimental settings, as the contribution from the weight changes dominates the contribution from the initialization. These findings validate that the C-NFC is an important contributor to increases in the UC-NFC.
Our experiments suggest that studying the C-NFC is a useful first step to understanding the neural feature ansatz. In the remainder of this section, we develop a theoretical analysis to understand why the C-NFC is generically large at early times. We then use this analysis to motivate interventions which can keep the C-NFC high and promote the NFA earlier on in training (Section 5).
3.1 Analysis at early times
We now theoretically identify why the centered AGOP and NFM become correlated as a structural property of gradient descent, for at least the early training times. We consider the setup introduced in Equation (1). For theoretical convenience we focus on the case of training under gradient flow, where the dynamics of a weight matrix $W^{(\ell)}$ trained on loss $\mathcal{L}$ is given by
$\dot{W}^{(\ell)}\equiv\frac{dW^{(\ell)}}{dt}=-\nabla_{W^{(\ell)}}\mathcal{L}~{},$ (4)
where $\dot{W}^{(\ell)}$ is the time derivative of $W^{(\ell)}$. We expect our results to apply to gradient descent with small and intermediate learning rates as well. For simplicity of notation, we omit the layer index $\ell$.
We note that $\bar{W}=0$. This gives us the following proposition, which shows that early time dynamics of the centered NFM and AGOP are dominated by the second time derivatives:
Proposition 3 (Centered NFC dynamics).
Let $W$ be the weights of a fully-connected layer of a neural network at initialization and $X$ be the inputs to that layer. Then, when the neural network is trained by gradient flow on a loss function $\mathcal{L}$, we have $\bar{W}^{\top}\bar{W}=\bar{W}^{\top}K\bar{W}=0$, $\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}\bar{W})=\frac{\operatorname{d}}{\operatorname{d}t}(\bar{W}^{\top}K\bar{W})=0$, and the first non-trivial time derivatives satisfy,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=2\cdot X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X,\quad\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})=2\cdot X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}$ (5)
@@ 3.1 Analysis at early times
where $\dot{\mathcal{L}}$ is the $n\times n$ diagonal matrix of logit derivatives ˙$\dot{\mathcal{L}}\equiv{{\rm diag}\left(\frac{\partial\mathcal{L}}{\partial f}\right)}$.
Proof of Proposition 3.
At initialization, all terms containing at least one copy of $\bar{W}$ vanish, leading to the zero and first time derivatives to be 0. Further, we have $\dot{W}=\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$. Therefore, using that $K\equiv\frac{\partial f(X)}{\partial h}^{\top}\frac{\partial f(X)}{\partial h}$ and $\mathcal{K}\equiv\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}$,
$\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=\dot{W}^{\top}\dot{W}=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X=X^{\top}\dot{\mathcal{L}}\mathcal{K}\dot{\mathcal{L}}X~{},$
and,
$\frac{1}{2}\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W})=\dot{W}^{\top}K\dot{W}$ $=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\,K\,\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\frac{\partial f(X)}{\partial h}\frac{\partial f(X)}{\partial h}^{\top}\dot{\mathcal{L}}X$
$=X^{\top}\dot{\mathcal{L}}\mathcal{K}^{2}\dot{\mathcal{L}}X~{}.$
We immediately see from these equations how gradient-based training can drive the C-NFC to large (close to $1$) initial values - the first non-trivial time derivatives are highly correlated. This also sheds light on another mystery: namely, that the NFC does not explicitly involve the loss, yet it holds after gradient based training on the loss. The C-NFC is driven by the fact that the dynamics of the centered NFM and AGOP depend on the labels in a similar way, differing only by a single factor of the kernel $\mathcal{K}$. This eventually drives the UC-NFC and causes the NFA to hold.
@@
In the remainder of this work, we first analyze the C-NFC in early training and show that gradient based training drives the C-NFC, with a strength that depends on model and datasets (Section 4). We then discuss initializations and optimizers which can increase the C-NFC and induce larger correlation between $F$ and $\bar{G}$ (Section 5).
4 Theoretically predicting the centered NFC
In this section, we show how our decomposition for the time derivatives of the centered NFC in early times can be used to predict its value theoretically.
4.1 Early time C-NFC dynamics
In order to understand the correlation of centered AGOP and NFM more quantitatively, we will focus on the case of mean-squared error (MSE) loss. Here $\dot{\mathcal{L}}$ corresponds to the diagonal matrix of the residuals $y-f(x)$. If the outputs of the network is 00 on the training data, then ˙$\dot{\mathcal{L}}=Y\equiv{\rm diag}\left(y\right)$, the labels themselves. In that case, the correlation of the time derivatives at initialization is a function of the input-label covariances as well as the PTK matrix:
@@ 4.1 Early time C-NFC dynamics
$\begin{split}&\mathbb{E}_{\theta}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right]=\\ &\text{tr}\left(X^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}\right]YXX^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}^{2}\right]YX\right)\\ &+\text{tr}\left(\text{Cov}\left(X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\right)\right)\end{split}$ (7)
A similar decomposition exists for the denominator terms.
The simplest limit is the NTK regime where the width tends to infinity with fixed number of data points and input dimension. Here the PTK matrix will approach its expectation, and the second, covariance term above will tend to 0. This suggests that the first (mean) term encodes much of the interesting phenomenology near initialization for large networks. We use this fact to design a dataset that interpolates between small and large initial derivatives of the C-NFC, even in the non-NTK regime, which will help us demonstrate the validity of our random matrix theory approach.
We focus most of our random matrix theory analysis on a one-hidden layer network, as the calculations quickly become complicated with depth. However, we also show that the mean term can be used to approximately predict the NFC value for Gaussian data with different spectral decay rates for more complicated networks (see Appendix D).
4.2 Exact predictions with one hidden layer and quadratic activations
@@
To illustrate our analysis, we derive here the numerator of Equation (6). With one hidden-layer, the neural network simplifies to $f(x)=a^{\top}(Wx)^{2}$, where $a$ is the readout layer, and the square is element-wise. Let $M^{(4)}_{X|Y}=(X^{\top}YX)^{2}$, and $M^{(2)}_{X}=X^{\top}X$, and $F_{a}=W^{\top}{\rm diag}\left(a^{2}\right)W$. Then,
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(M^{(4)}_{X|Y}F_{a}M^{(4)}_{X|Y}F_{a}M^{(2)}_{X}F_{a}\right)~{}.$
@@
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(M^{(4)}_{X|Y}\right)^{2}\text{tr}\left(F_{a}^{3}\right)+\text{tr}\left(F_{a}\right)\text{tr}\left(F_{a}^{2}\right)\text{tr}\left((\bar{M}^{(4)}_{X|Y})^{2}\right)~{}.$
@@
To numerically explore the validity of the random matrix theory calculations, we need a way to generate datasets with different values of $\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)$. We construct a random dataset called the alignment reversing dataset, parameterized by a balance parameter $\gamma\in(0,1]$ to adversarially disrupt the NFA near initialization in the regime that width $k$, input dimension $d$, and dataset size $n$ are all equal ($n=k=d=1024$). By Proposition 8, for the aforementioned neural architecture, the expected second derivative of the centered NFM satisfies, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YX=(X^{\top}YX)^{2}$, while the expected second derivative of the centered AGOP, $\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX$, has an additional component $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$. Our construction exploits this difference in that $X^{\top}X$ becomes adversarially unaligned to $X^{\top}YX$ as the balance parameter decreases.
@@
The construction exploits that we can manipulate $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$ freely of the NFM using a certain choice of $Y$. We design the dataset such that this AGOP-unique term is close to identity, while the NFM second derivative has many large off-diagonal entries, leading to low correlation between the second derivatives of the NFM and AGOP.
Dataset Construction
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by a balance parameter $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
$\Sigma=\bm{1}\bm{1}^{\top}+\epsilon_{1}\cdot I~{}.$
Then the second subset, $X^{(2)}$, is constructed such that $(X^{(2)})^{\top}X^{(2)}\approx((X^{(1)})^{\top}X^{(1)})^{-2}$, and has labels $y_{2}=0$. Then, for balance parameter $\gamma$ sufficiently small, the AGOP second derivative approximately satisfies,
$=(X^{(1)})^{\top}X^{(1)}X^{\top}X(X^{(1)})^{\top}X^{(1)}$
$\approx(X^{(1)})^{\top}X^{(1)}(X^{(2)})^{\top}X^{(2)}(X^{(1)})^{\top}X^{(1)}$
$\approx I~{},$
In contrast, the NFM second derivative, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=(X^{\top}YX)^{2}=((X^{(1)})^{\top}X^{(1)})^{2}\approx\Sigma^{2}$, will be significantly far from identity.
In our experiment, we sample multiple random datasets with this construction and compute the predicted and observed correlation of the second derivatives of the centered NFC at initialization. For specific details on the construction, see Appendix F.
@@
We observe in Figure 3 that the centered NFC predicted with random matrix theory closely matches the observed values, across individual four random seeds and for the average of the correlation across them. Crucially, a single neural network is used across the datasets, confirming the validity of the self-averaging assumption. The variation in the plot across seeds come from randomness in the sample of the data, which cause deviations from the adversarial construction.
5 Increasing the centered contribution to the NFC strengthens feature learning